You decide to invest in gold: you purchase 30 ounces of gold at $1,550 per ounce and hold it for two years. At the end of the two years you sell your gold for $1,700 per ounce. What is the return on your investment?
Answer #1

voces (fw = pw (Hot Where; fwa future werth Pw=present worth tatire r= return (info) 1700 - 1550 (148)2 1.096 = (1+2 (ltos a